# World Models Increase Autonomy in Reinforcement Learning

### v2

@@ Abstract
Reinforcement learning (RL) is an appealing paradigm for training intelligent agents, enabling policy acquisition from the agent’s own autonomously acquired experience. However, the training process of RL is far from automatic, requiring extensive human effort to reset the agent and environments. To tackle the challenging reset-free setting, we first demonstrate the superiority of model-based (MB) RL methods in such setting, showing that a straightforward adaptation of MBRL can outperform all the prior state-of-the-art methods while requiring less supervision. We then identify limitations inherent to this direct extension and propose a solution called model-based reset-free (MoReFree) agent, which further enhances the performance. MoReFree adapts two key mechanisms, exploration and policy learning, to handle reset-free tasks by prioritizing task-relevant states. It exhibits superior data-efficiency across various reset-free tasks without access to environmental reward or demonstrations while significantly outperforming privileged baselines that require supervision. Our findings suggest model-based methods hold significant promise for reducing human effort in RL. Website: https://sites.google.com/view/morefree
@@ 3.2 Model-based RL setup
Recent goal-conditioned MBRL approaches like LEXA (Mendonca et al., 2021) and PEG (Hu et al., 2023) train goal-conditioned policies purely using synthetic data generated by learned world models. Their robust exploration demonstrate significant success in solving long-horizon goal-conditioned tasks. In the reset-free setting, strong exploration is crucial, as the agent can no longer depend on episodic resets to bring it back to task-relevant areas if it gets stuck. Therefore, we select PEG as the backbone MBRL agent for its strong exploration abilities and sample efficiency.
@@
The world model is a recurrent state-space model (RSSM) which is trained to predict future states and is used as a learned simulator to train the policies and value functions. The goal-conditioned policy is trained to reach random states sampled from the replay buffer. The exploration policy is trained on an intrinsic motivation reward that rewards world model error, expressed through the variance of an ensemble (Sekar et al., 2020). Both policies are trained on simulated trajectory rollouts in the world model.
Self-supervised Goal-reaching Reward Function: Rather than assuming access to the environmental reward, PEG learns its own reward function. PEG uses a dynamical distance function (Hartikainen et al., 2019) as the reward function within world models, which predicts the number of actions between a start and goal state. The distance function is trained on random state pairs from imaginary rollouts of . is then trained to minimize the dynamical distance between its states and commanded goal state in imagination. See Mendonca et al. (2021) for more details.
@@
Then, in the “Explore”-phase, starting from the last state in the “Go”-phase, the exploration policy is run for steps, resulting in . The interleaving of goal-conditioned behavior with exploratory behavior results in more directed exploration and informative data. This in turn improves accuracy of the world model, and the policies that train inside the world model. See Algorithm 1 and Algorithm 2 for pseudocode. The choice of goal distribution is important for Go-Explore. In easier tasks, the evaluation goal distribution may be sufficient. But in longer-horizon tasks, evaluation goals may be too hard to achieve. Instead, intermediate goals from an exploratory goal distribution can help the agent explore. We choose PEG, which generates goals by planning through the world model to maximize exploration value (see Hu et al. (2023) for details).
@@ 4.1 Back-and-Forth Go-Explore
Now, we define the “Back-and-Forth Go-Explore” strategy as seen in Algorithm 3. First, we decide whether to
perform initial / evaluation state directed exploration. With probability , we sample goals from respectively. Then, we execute the Go-Explore routine for each goal. We name Go-Explore trajectories conditioned on initial state goals as “Back” trajectories, and Go-Explore trajectories conditioned on evaluation goals as “Forward” trajectories. With probability , we execute exploratory Go-Explore behavior by sampling exploratory goals from PEG. The difference between reset-free PEG and MoReFree can be seen in Algorithm 2, unlike PEG, MoReFree employs the “Back-and-Forth Go-Explore”.
@@ 4.3 Implementation Details
Our work builds on the top of PEG (Hu et al., 2023), and we use its default hyperparameters for world model, policies, value functions and temporal reward function. We set the length of each phase for Go-Explore to half the evaluation episode length for each task. We set the default value of for all tasks (never tuned). See Section C.3 for more details and the supplemental for MoReFree code.
@@ 5 Experiments
Baselines: All baselines except for R3L are implemented using official codebases, see Section C.2 for details.
@@
Environments: We evaluate MoReFree and baselines on eight tasks (see Figure 3). We select five tasks from IBC’s evaluation suite of six tasks; (Fetch Reach is omitted because it is trivially solvable). Next, we increased the complexity of the two hardest tasks from IBC, Fetch Push and Fetch Pick&Place, by extending the size of the workspace, replacing artificial workspace limits (which cause unrealistic jittering behavior near the limits, see the website for videos) with real walls, and evaluating on harder goal states (i.e. Pick&Place goals only in the air rather than including ones on the ground). In addition, we contributed a difficult locomotion task, Ant, which is adapted from the PEG codebase (Hu et al., 2023). All methods are run with 5 seeds, and the mean performance and standard error are reported. During the evaluation, the performance on tasks with randomly sampled goals from is measured by averaging over 10 episodes. See Appendix C for more experimental details.
@@ 5.3 Ablations
To justify our design choices, we ablate the two mechanisms of MoReFree, the back-and-forth exploration and task-relevant goal-conditioned policy training, and plot the results in Figure 7 First, removing all mechanisms (MF w/o Explore & Imag.) reduces to reset-free PEG, and we can see a large gap in performance. Next, MF with Only Task Goals sets , which causes an extreme bias towards task-relevant states in the exploration and policy training. This also degrades performance, due to the need for strong exploration in the reset-free setting. Examinations of more values for can be found in Section C.3.
@@ 6 Conclusion and Future Work
As a step towards reset-free training, we adapt model-based methods to the reset-free setting and demonstrate their superior performance. Specifically, we show that with minor modifications, unsupervised MBRL method substantially outperforms the state-of-the-art model-free baselines tailored for the reset-free setting while being more autonomous (requires less supervision like environmental reward or demonstrations). We then identify a limitation of unsupervised MBRL in the reset-free setting (over-exploration on task-irrelevant states), and propose MoReFree to address such limitations by focusing model-based exploration and goal-conditioned policy training on task-relevant states. We conduct a through experimental study of MoReFree and baselines over 8 tasks, and show considerable performance gains over the MBRL baseline and prior state-of-the-art reset-free methods. Despite its overall success, MoReFree is not without limitations. Being a model-based approach, it inherits all associated disadvantages. For example, we believe Sawyer Door is a task where learning the dynamics is harder than learning the policy (see Appendix F), disadvantaging MBRL approaches. Next, MoReFree uses a fixed percentage of task-relevant goals for exploration and imagination, whereas future work could consider an adaptive curriculum. Finally, scaling MoReFree to high-dimensional observations would be a natural extension. We hope MoReFree inspires future efforts in increasing autonomy in RL.
@@ Appendix A Broader Impacts
Appendix B Extended Related Work
Learned Reward Functions: Instead of requiring the environment to provide a reward function, the agent can learn its own reward function from onboard sensors and data. Given human specified example states, e.g. a goal image, VICE and C-Learning train reward classifiers over examples (Fu et al., 2018; Eysenbach et al., 2021) and agent data. The learned dynamical distance function (Hartikainen et al., 2019) learns to predict the number of actions between pairs of states. The dynamical distance function is used by unsupervised MBRL approaches like LEXA and PEG (Mendonca et al., 2021; Hu et al., 2023) to train the goal-conditioned policy.
Appendix C Experimental Details
C.1 Environments
@@
Push (hard): Different from the original Fetch Push task, in our case walls are added to prevent the block from dropping out of the table. The workspace of the robot arm is also limited. The block is always initialized to a fixed location, and goal distribution during the evaluation is . Fetch Push used in the IBC paper, the block is limited by joint constraint, which shows unrealistic jittering behaviors near the limits (we observe such phenomenon by running model-based go-explore, the exploration policy prefers to always interact with the block and keep pushing it towards the limit boundary, see videos on our project website 111https://sites.google.com/view/morefree). Meanwhile, the gripper is blocked, which makes the task easier. In our case, we release the gripper and it can now open and close again which add two more dimension of the state space. We found it is important to release the gripper in our version of Push task, when the block is in corners, it will need to operate the gripper to drag the block escape from corners. The maximum steps the agent can take in 50 during the evaluation. Hard reset will happen after every steps. In the whole training process we performed, it resets 5 times in total.

### v3

# Reset-free Reinforcement Learning with World Models

@@ Abstract
Reinforcement learning (RL) is an appealing paradigm for training intelligent agents, enabling policy acquisition from the agent’s own autonomously acquired experience. However, the training process of RL is far from automatic, requiring extensive human effort to reset the agent and environments. To tackle the challenging reset-free setting, we first demonstrate the superiority of model-based (MB) RL methods in such setting, showing that a straightforward adaptation of MBRL can outperform all the prior state-of-the-art methods while requiring less supervision. We then identify limitations inherent to this direct extension and propose a solution called model-based reset-free (MoReFree) agent, which further enhances the performance. MoReFree adapts two key mechanisms, exploration and policy learning, to handle reset-free tasks by prioritizing task-relevant states. It exhibits superior data-efficiency across various reset-free tasks without access to environmental reward or demonstrations while significantly outperforming privileged baselines that require supervision. Our findings suggest model-based methods hold significant promise for reducing human effort in RL. Website: https://yangzhao-666.github.io/morefree
@@ 2 Related Work
Learned Reward Functions: Instead of requiring the environment to provide a reward function, the agent can learn its own reward function from onboard sensors and data. Given human specified example states, e.g. a goal image, VICE and C-Learning train reward classifiers over examples (Fu et al., 2018; Eysenbach et al., 2021) and agent data. The learned dynamical distance function (Hartikainen et al., 2019) learns to predict the number of actions between pairs of states. The dynamical distance function is used by unsupervised MBRL approaches like LEXA and PEG (Mendonca et al., 2021; Hu et al., 2023) to train the goal-conditioned policy. MoReFree also employs the dynamical distance function as the reward function to eliminate the need of the environmental reward.
@@ 3.2 Model-based RL setup
Recent goal-conditioned MBRL approaches like LEXA (Mendonca et al., 2021) and PEG (Hu et al., 2023) train goal-conditioned policies purely using synthetic data generated by learned world models. Their robust exploration demonstrates significant success in solving long-horizon goal-conditioned tasks. In the reset-free setting, strong exploration is crucial, as the agent can no longer depend on episodic resets to bring it back to task-relevant areas if it gets stuck. Therefore, we select PEG as the backbone MBRL agent for its strong exploration abilities and sample efficiency.
@@
The world model is a recurrent state-space model (RSSM) which is trained to predict future states and is used as a learned simulator to train the policies and value functions. The goal-conditioned policy is trained to reach random states sampled from the replay buffer. The exploration policy is trained on an intrinsic motivation reward that rewards world model error, expressed through the variance of an ensemble (Sekar et al., 2020)(see Section C.3 for more details). Both policies are trained on simulated trajectory rollouts in the world model.
Self-supervised Goal-reaching Reward Function: Rather than assuming access to the environmental reward, PEG learns its own reward function. PEG uses a dynamical distance function (Hartikainen et al., 2019) as the reward function within world models, which predicts the number of actions between a start and goal state. The distance function is trained on random state pairs from imaginary rollouts of . is then trained to minimize the dynamical distance between its states and commanded goal state in imagination. See Section C.1 for more details.
@@
Then, in the “Explore”-phase, starting from the last state in the “Go”-phase, the exploration policy is run for steps, resulting in . The interleaving of goal-conditioned behavior with exploratory behavior results in more directed exploration and informative data. This in turn improves accuracy of the world model, and the policies that train inside the world model. See Algorithm 1 and Algorithm 2 for pseudocode. The choice of goal distribution is important for Go-Explore. In easier tasks, the evaluation goal distribution may be sufficient. But in longer-horizon tasks, evaluation goals may be too hard to achieve. Instead, intermediate goals from an exploratory goal distribution can help the agent explore. We choose PEG, which generates goals by planning through the world model to maximize exploration value (see Section C.2 for details).
@@ 4.1 Back-and-Forth Go-Explore
Now, we define the “Back-and-Forth Go-Explore” strategy as seen in Algorithm 3. First, we decide whether to perform initial / evaluation state directed exploration. With probability , we sample goals from respectively. Then, we execute the Go-Explore routine for each goal. We name Go-Explore trajectories conditioned on initial state goals as “Back” trajectories, and Go-Explore trajectories conditioned on evaluation goals as “Forward” trajectories. With probability , we execute exploratory Go-Explore behavior by sampling exploratory goals from PEG. The difference between reset-free PEG and MoReFree can be seen in Algorithm 2, unlike PEG, MoReFree employs the “Back-and-Forth Go-Explore”.
@@ 4.3 Implementation Details
Our work builds on the top of PEG (Hu et al., 2023), and we use its default hyperparameters for world model, policies, value functions and temporal reward function. We set the length of each phase for Go-Explore to half the evaluation episode length for each task. We set the default value of for all tasks (never tuned). See Section B.3 for more details and the supplemental for MoReFree code.
@@ 5 Experiments
Baselines: All baselines except for R3L are implemented using official codebases, see Section B.2 for details.
@@
DreamerV2 (Hafner et al., 2020) is a commonly used MBRL method. The goal-conditioned policy is executed for the whole reset-free episode.
@@
Environments: We evaluate MoReFree and baselines on eight tasks (see Figure 3). We select five tasks from IBC’s evaluation suite (Kim et al., 2023) of six tasks; (PointUMaze, Tabletop, Sawyer Door, Fetch Push and PP, Fetch Reach is omitted because it is trivially solvable). Next, we increased the complexity of the two hardest tasks from IBC, Fetch Push and Fetch Pick&Place, by extending the size of the workspace, replacing artificial workspace limits (which cause unrealistic jittering behavior near the limits, see the website for videos) with real walls, and evaluating on harder goal states (i.e. Pick&Place goals only in the air rather than including ones on the ground). In addition, we contributed a difficult locomotion task, Ant, which is adapted from the PEG codebase (Hu et al., 2023).
PointUMaze: A point-mass agent navigates a U-shape maze through continuous acceleration commands. During evaluation, the agent starts from the bottom-left corner and is tasked to reach the top-left corner.
Tabletop Manipulation: The agent needs to grab and move the mug to one of the four goal locations. The initial state is always fixed and the goal state is uniformly sampled from four fixed locations.
Sawyer Door: The agent controls a Sawyer robot arm to close the door in an open position. During the reset-free training, it needs to learn to close the door and open the door again to practice. The door is opened to 60 degrees for evaluation.
Fetch Push&PP: The agent commands a Fetch robot arm to push / pick&place the object initialized at the center of the table to goal locations. The environment is taken from IBC’s evaluation suite, which modified the original environment from Plappert et al. (2018). To prevent the block from falling off the table, the IBC authors artificially limited the block position with block position constraints.
Push&PP(hard): Using block position constraints (in Fetch Push&PP) resulted in unrealistic jittering behavior near the limits. To avoid this, we removed the artificial joint constraints and surrounded the table with physical walls. Furthermore, we enable the usage of the grippers (disabled in IBC’s version) to permit picking behaviors (i.e. useful for resetting), at the cost of increased action space and exploration difficulty.
Ant: The 4-legged ant agent needs to navigate in a square room to a given goal, which is uniformly located in the top-right corner. The initial state is at the center point with randomness. It is adapted from Hu et al. (2023), with changing the U-shape maze into a square room.
Most methods are run with 5 seeds, and the mean performance and standard error are reported. During the evaluation, the performance on tasks with randomly sampled goals from is measured by averaging over 10 episodes. See Appendix B for more experimental details.
@@ 5.3 Ablations
To justify our design choices, we ablate the two mechanisms of MoReFree, the back-and-forth exploration and task-relevant goal-conditioned policy training, and plot the results in Figure 7.
First, removing all mechanisms (MF w/o Explore & Imag.) reduces to reset-free PEG, and we can see a large gap in performance. Next, MF with Only Task Goals sets , which causes an extreme bias towards task-relevant states in the exploration and policy training. This also degrades performance, due to the need for strong exploration in the reset-free setting. Examinations of more values for can be found in Section B.3.
@@ 6 Conclusion and Future Work
As a step towards reset-free training, we adapt model-based methods to the reset-free setting and demonstrate their superior performance. Specifically, we show that with minor modifications, unsupervised MBRL method substantially outperforms the state-of-the-art model-free baselines tailored for the reset-free setting while being more autonomous (requires less supervision like environmental reward or demonstrations). We then identify a limitation of unsupervised MBRL in the reset-free setting (over-exploration on task-irrelevant states), and propose MoReFree to address such limitations by focusing model-based exploration and goal-conditioned policy training on task-relevant states. We conduct a thorough experimental study of MoReFree and baselines over 8 tasks, and show considerable performance gains over the MBRL baseline and prior state-of-the-art reset-free methods.
Despite its overall success, MoReFree is not without limitations. Being a model-based approach, it inherits all associated disadvantages. For example, we believe Sawyer Door is a task where learning the dynamics is harder than learning the policy (see Appendix F), disadvantaging MBRL approaches. Next, MoReFree uses a fixed percentage of task-relevant goals for exploration and imagination, whereas future work could consider an adaptive curriculum. Finally, scaling MoReFree to high-dimensional observations and real-world applications would be natural extensions. We hope MoReFree inspires future efforts in increasing autonomy in RL.
@@ Appendix A Broader Impacts
Appendix B Experimental Details
B.1 Environments
@@
Push (hard): Different from the original Fetch Push task, in our case walls are added to prevent the block from dropping out of the table. The workspace of the robot arm is also limited. The block is always initialized to a fixed location, and goal distribution during the evaluation is . Fetch Push used in the IBC paper, the block is limited by joint constraint, which shows unrealistic jittering behaviors near the limits (we observe such phenomenon by running model-based go-explore, the exploration policy prefers to always interact with the block and keep pushing it towards the limit boundary, see videos on our project website 111https://yangzhao-666.github.io/morefree). Meanwhile, the gripper is blocked, which makes the task easier. In our case, we release the gripper and it can now open and close again which add two more dimension of the state space. We found it is important to release the gripper in our version of Push task, when the block is in corners, it will need to operate the gripper to drag the block escape from corners. The maximum steps the agent can take in 50 during the evaluation. Hard reset will happen after every steps. In the whole training process we performed, it resets 5 times in total.
@@ B.2 Baseline Implementations
DreamerV2: We use the official implementation of PEG. In order to reduce it to DreamerV2 (Hafner et al., 2020), we remove the exploration policy and only execute goal-conditioned policy for the whole reset-free episode. During imagination training, the goal-conditioned policy is only trained on the evaluation goal distribution.
@@ B.5 Resource Usage
Appendix C Method Details
Here, we provide a more in-depth exposition of self-supervised goal-reaching reward function (Hartikainen et al., 2019) that is used for training the goal-conditioned policy, PEG (Hu et al., 2023) that is used for generating exploratory goal distribution, and P2E (Sekar et al., 2020) that is used for training the exploration policy.
C.1 Self-supervised Goal-reaching Reward Function
MoReFree does not require environment reward functions, instead it learns a distance function for training the goal-conditioned policy. is trained by sampling pairs of states from an imagined rollout of the goal-conditioned policy and predicting the distance , where is the maximum distance equal to the imagination horizon. Then the reward is defined as .
C.2 Exploratory Goal Distribution
We use PEG to generate the exploratory goal distribution in MoReFree. PEG generates goals that have high exploration potentials. To evaluate a goal , the goal-conditioned policy is rolled out for trajectories within the learned world model. Then the terminal state exploration value for each trajectory with the learned exploration value function is estimated, where is the last state of the trajectory . Then the estimates are averaged.
The goal variable is optimized using model predictive path integral control (MPPI). First, goal candidates are sampled from an initial distribution. These candidates are then evaluated as described above. This averaged exploration value acts as the “score" for the goal candidate. Once we have scores for each goal candidate, a Gaussian distribution is fit according to the rule:
where is the reward weight hyperparameter. We then sample candidates from the computed Gaussian, and repeat the process for multiple iterations. After the last iteration, is defined as the computed Gaussian.
C.3 Plan2Explore
The world model consists of the following components:
The model states contain a deterministic component and a stochastic component .
P2E is the objective we used to train the exploration policy, and it encourages the agent to visit states that can improve the world model the most. We train an ensemble of 1-step models to predict the next model state from the current model state:
Then the exploration reward is the variance of the ensemble predictions averaged across dimension of the model state, .